OpenStudy (calculusxy):

Find the tenth term in the sequence. \(a_n = na_{n-1}\) \(a_1 = -1\)

OpenStudy (calculusxy):

@jim_thompson5910

myininaya (myininaya):

can you find a2 first? there might be a pattern we can find so we don't have to go all the way to a10

myininaya (myininaya):

to find a2 replace the n's with 2

OpenStudy (calculusxy):

Maybe.. \(a_2 = 2a_{2-1}\) \(a_2 = 2(a_1)\) \(a_2 = 2(-1)\) \(a_2 = -2\)

myininaya (myininaya):

that's right

OpenStudy (calculusxy):

So the common ratio is 2?

myininaya (myininaya):

no this isn't a geometric sequence

OpenStudy (calculusxy):

Why?

myininaya (myininaya):

geometric sequence would be a_n=constant*a_(n-1) and here the common ration would be that constant but n is not a constant it changes

OpenStudy (calculusxy):

I learned the geometric sequence formula is like: \(a_n = a_1(r)^{n-1}\). Is it the same thing as what you wrote?

myininaya (myininaya):

the thing you wrote is an explicit form of a geometric sequence the thing I wrote is implicit form of a geometric sequence

OpenStudy (calculusxy):

I didn't learn about the implicit form.

myininaya (myininaya):

\[a_n=a_1 r^{n-1} \\ \text{ then } a_{n-1}=a_1 r^{(n-1)-1}= a_1 r^{n-2} \\ \text{ and } \frac{a_n}{a_{n-1}}=\frac{a_1 r^{n-1}}{a_1 r^{n-2}} \\ \implies \frac{a_n}{a_{n-1}}=\frac{r^{n-1}}{r^{n-2}} \\ \implies \frac{a_n}{a_{n-1}}=r^{(n-1)-(n-2)} \\ \implies \frac{a_n}{a_{n-1}}= r^{1} \\ \\ \implies a_n= r \cdot a_{n-1}\]

myininaya (myininaya):

I started with your explicit form for a geometric sequence and I found a_(n-1) by replacing n with n-1 and then the steps after is on the way to revealing the implicit form like I divide the a_n thing by the a_(n-1) thing to show our equations were equivalent

OpenStudy (calculusxy):

OK thank you for your help :)

myininaya (myininaya):

no problem :)